Consider the following systemof linear equations:



x+2y+2z=1,x+ay+3z=3,x+11y+az=b.




For which values of a does the system have a unique solution. and for which pairs of values (a,b) does the system have more than on solution?

1
Expert's answer
2022-02-10T17:38:29-0500

"x+2y+2z=1""x+ay+3z=3""x+11y+az=b"

Augmented matrix


"\begin{bmatrix}
 1 & 2 & 2 && 1 \\
 1 & a & 3 && 3 \\
 1 & 11 & a && b \\
\end{bmatrix}"

"R_2=R_2-R_1"


"\begin{bmatrix}
 1 & 2 & 2 && 1 \\
 0 & a-2 & 1 && 2 \\
 1 & 11 & a && b \\
\end{bmatrix}"

"R_3=R_3-R_1"


"\begin{bmatrix}
 1 & 2 & 2 && 1 \\
 0 & a-2 & 1 && 2 \\
 0 & 9 & a -2&& b-1 \\
\end{bmatrix}"

If "a-2=0"


"\begin{bmatrix}
 1 & 2 & 2 && 1 \\
 0 & 0 & 1 && 2 \\
 0 & 9 & 0&& b-1 \\
\end{bmatrix}"

"y=\dfrac{b-1}{9}, z=2, x=1-\dfrac{2(b-1)}{9}-4"

If "a=2," we have the unique solution


"(-\dfrac{2b+25}{9}, \dfrac{b-1}{9}, 2)"

If "a\not=2"

"R_2=\dfrac{R_2}{a-2}"


"\begin{bmatrix}
 1 & 2 & 2 && 1 \\
\\
 0 & 1 & \dfrac{1}{a-2} && \dfrac{2}{a-2} \\
\\
 0 & 9 & a -2&& b-1 \\ 
\end{bmatrix}"

"R_1=R_1-2R_2"


"\begin{bmatrix}
 1 & 0 & 2- \dfrac{2}{a-2} && 1-\dfrac{4}{a-2} \\
\\
 0 & 1 & \dfrac{1}{a-2} && \dfrac{2}{a-2} \\
\\
 0 & 9 & a -2&& b-1 \\ 
\end{bmatrix}"

"R_3=R_3-9R_2"


"\begin{bmatrix}
 1 & 0 & \dfrac{2a-6}{a-2} && \dfrac{a-6}{a-2} \\
\\
 0 & 1 & \dfrac{1}{a-2} && \dfrac{2}{a-2} \\
\\
 0 & 0 & \dfrac{(a-2)^2-9}{a-2} && \dfrac{(b-1)(a-2)-18}{a-2} \\ 
\end{bmatrix}"

If "\dfrac{(a-2)^2-9}{a-2}=0"


"a_1=-1, a_2=5"

"a=-1"


"\begin{bmatrix}
 1 & 0 & 8/3 && 7/3 \\

 0 & 1 & -1/3 && -2/3 \\

 0 & 0 & 0 && b+5 \\ 
\end{bmatrix}"

If "b\not=-5," the system has no solution.

If "b=-5," the system has more than one solution.


"a=5"


"\begin{bmatrix}
 1 & 0 & 4/3 && -1/3 \\

 0 & 1 & 1/3 && 2/3 \\

 0 & 0 & 0 && b-7 \\ 
\end{bmatrix}"

If "b\not=7," the system has no solution.

If "b=7," the system has more than one solution.


If "a\not=2, a\not=-1, a\not=5"

"R_3=(\dfrac{a-2}{(a-2)^2-9})R_3"


"\begin{bmatrix}
 1 & 0 & \dfrac{2a-6}{a-2} && \dfrac{a-6}{a-2} \\
\\
 0 & 1 & \dfrac{1}{a-2} && \dfrac{2}{a-2} \\
\\
 0 & 0 & 1 && \dfrac{(b-1)(a-2)-18}{(a-2)^2-9} \\ 
\end{bmatrix}"

"R_1=R_1-(\dfrac{2a-6}{a-2})R_3"


"\begin{bmatrix}
 1 & 0 & 0 && \dfrac{a^2-6a+33-2(a-3)b}{(a+1)(a-5)} \\
\\
 0 & 1 & \dfrac{1}{a-2} && \dfrac{2}{a-2} \\
\\
 0 & 0 & 1 && \dfrac{(b-1)(a-2)-18}{(a-2)^2-9} \\ 
\end{bmatrix}"

"R_2=R_2-\dfrac{R_3}{a-2}"


"\begin{bmatrix}
 1 & 0 & 0 && \dfrac{a^2-6a+33-2(a-3)b}{(a+1)(a-5)} \\
\\
 0 & 1 & 0 && \dfrac{2a-3-b}{(a+1)(a-5)} \\
\\
 0 & 0 & 1 && \dfrac{(b-1)(a-2)-18}{(a+1)(a-5)} \\ 
\end{bmatrix}"

In this case the system has the unique solution.


i) The system has the unique solution for "a\not=-1, a\not=5."


ii) The system has more than one solution for "(a, b)=(-1, -5)" or "(a, b)=(5, 7)."